# CRITERIA: a New Benchmarking Paradigm for Evaluating Trajectory Prediction Models for Autonomous Driving

### v1

@@ Abstract
In this paper, we propose a new benChmarking paRadIgm for evaluaTing trajEctoRy predIction Approaches (CRITERIA). Particularly, we propose 1) a method for extracting driving scenarios at varying levels of specificity according to the structure of the roads, models’ performance, and data properties for fine-grained ranking of prediction models; 2) A set of new bias-free metrics for measuring diversity, by incorporating the characteristics of a given scenario, and admissibility, by considering the structure of roads and kinematic compliancy, motivated by real-world driving constraints. 3) Using the proposed benchmark, we conduct extensive experimentation on a representative set of the prediction models using the large scale Argoverse dataset. We show that the proposed benchmark can produce a more accurate ranking of the models and serve as a means of characterizing their behavior. We further present ablation studies to highlight contributions of different elements that are used to compute the proposed metrics.
@@ I Introduction
Existing trajectory prediction benchmarks [1] [2] [3] consist of driving data and distance-based metrics to evaluate the accuracy of the models against the ground truth. The datasets, however, are typically biased towards more common scenarios, such as cruising. Consequently, when metrics are averaged over the entire data, the benchmark tends to favor the models that perform better on common cases and provides little insight into the differences among the models and how they behave in rarer situations. This would limit the ability to observe the effects of design choices on particular driving scenarios.
@@
Furthermore, we propose a novel set of biased-free metrics for accurately measuring the performance of trajectory prediction models. Our metrics present a new insight into the diversity and admissibility of trajectories by defining road and kinematic compliancy for normal driving behavior, which are vital aspects of real-world driving. Our metrics help characterize models’ behavior more effectively and provide a different, yet more accurate ranking scheme than the existing metrics.
@@ II-A Trajectory Prediction
An extensive body of literature is dedicated to trajectory prediction in the context of autonomous driving [7, 8, 9, 10, 11, 12, 13, 14, 15]. Existing methods investigate various representations and approaches to learning scene context, including rasterized images [16, 17, 18], point-clouds [19] processed with convolutional neural networks, and vectorized representations processed using different architectures, such as graph neural networks [20, 21, 22, 23] or transformer architectures[24, 25, 26, 27] often combined with sophisticated fusion mechanisms [28]. To address the inherent uncertainty in predicting future trajectories, many models resort to generating multimodal trajectories in order to cover the space of possibilities. There are different approaches to achieving this goal, including the use of generative adversarial networks [29] [30], conditional variational autoencoders [31], or training sampling networks [32, 5]. Some methods also use trajectory endpoints [22] [33] or targets [34, 22, 35, 33] to model possible future intentions of vehicles to encourage generation of a more diverse set of trajectories. Regardless of the method of choice, there is a need for an effective mechanism to evaluate the quality of generated trajectories both in terms of diversity and admissibility, i.e. compliancy to road structure or dynamical limitations.
@@ II-B1 Distance-based Metrics
II-B2 Admissibility Merics
Admissibility metrics determine whether trajectories satisfy map constraints, e.g. road boundaries. Some of the metrics in this category include Driveable Area Compliance (DAC) and off-road rate [1], which measure the ratio of trajectories that fall within the driveable area. Both of these metrics, however, consider the entire scene as the drivable area and do not penalize the model if the predictions go outside of the valid lanes, e.g. the ones in the opposite direction. Another metric is referred to as Driveable Area Occupancy (DAO) [4], which measures the proportion of pixels occupied by the trajectories in the drivable area. Similar to other metrics, DAO is highly biased towards the length of trajectories as well it does not constrain the extent of the drivable area. This means, the longer trajectories get the higher (better) the value of the metric.
@@ III-A Problem Formulation
Trajectory prediction task can be viewed as an optimization problem. At time step $t$, let consider the past trajectory of the $i$-th agent as a set of 2D coordinates in bird’s eye view over some observation horizon $L$ time steps $X_{i}=\{(x_{i},y_{i})^{t-L+1},\cdots,(x_{i},y_{i})^{t-1}\}$. Then, the objective is to predict future trajectories $Y_{i}=\{(x_{i},y_{i})^{t+1},\cdots,(x_{i},y_{i})^{t+T}\}$, where $T$ is the prediction horizon. In this formulation, additional contextual information, such as road lane boundaries and links are used. In the multimodal setting, prediction models output $K$ different trajectories (referred to as modes) per each vehicle. For simplicity, in the remainder of this paper, we refer to $t-L+1$ and $t+T$ as $L$ and $T$, respectively.
@@ III-B Scenario Extraction
Trajectory prediction models have different architectural designs and often behave differently when applied to the same samples. For instance, the same models may generate different distributions, trajectories with different lengths, curvature, etc. As a result, it is essential to extract driving scenarios in a meaningful way to better reflect such differences among the models. To this end, we propose a method to extract driving scenarios based on three criteria, namely the road structure, models’ performance, and data properties.
@@ III-C2 Road Boundary Alignment Test
$\mathcal{C}=max(0,1-\frac{\Delta\Theta}{\pi})$ (1)
@@ III-D2 AMV: Average Magnitude Variation
where $\|\vec{v^{\prime}}^{t}\|$ is the magnitude of the trajectory vector $\vec{v}=\langle\hat{x}^{t}-\hat{x}^{t-1},\hat{y}^{t}-\hat{y}^{t-1}\rangle$ starting at time step $t-1$ and ending at $t$ that is passed by kinematic compliancy test. The magnitude variation between two trajectories is calculated once $MV_{ij}=MV_{ji}$, and then AMV is calculated by taking an average over all trajectory pairs’ magnitude variances.
@@ IV-B Accuracy, Diversity, and Admissibility
The proposed diversity metrics, show an alternative ranking. Thanks to removing length bias, we can see that MMTransformer achieves the highest magnitude diversity. This new ranking suggests that TNT tends to generate longer trajectories, compared to other models, hence being favored in diversity metrics in general. This model, however, achieves a second best score in terms of angular diversity owing to target selection based on road topology.
@@
Figure 2: Performance comparison on the challenging scenarios in terms of diversity, admissibility, and minFDE (represented as the size of circles, so the smaller is better).
@@ IV-C Scenario-based Analysis
Figure 3: Qualitative samples of different ATT test scores for predicted trajectories. The red lines indicate inadmissible trajectories. From left to right, the samples are generated by TNT, LaneGCN, TNT and MMTransformer.

### v2

@@ Abstract
In this paper, we propose a new benChmarking paRadIgm for evaluaTing trajEctoRy predIction Approaches (CRITERIA). Particularly, we propose 1) a method for extracting driving scenarios at varying levels of specificity according to the structure of the roads, models’ performance, and data properties for fine-grained ranking of prediction models; 2) A set of new bias-free metrics for measuring diversity, by incorporating the characteristics of a given scenario, and admissibility, by considering the structure of roads and kinematic compliancy, motivated by real-world driving constraints; 3) Using the proposed benchmark, we conduct extensive experimentation on a representative set of the prediction models using the large scale Argoverse dataset. We show that the proposed benchmark can produce a more accurate ranking of the models and serve as a means of characterizing their behavior. We further present ablation studies to highlight contributions of different elements that are used to compute the proposed metrics [^1] Code available at https://github.com/huawei-noah/SMARTS/tree/CRITERIA-latest/papers/CRITERIA.
@@ I Introduction
Existing trajectory prediction benchmarks [1, 2, 3] consist of driving data and distance-based metrics to evaluate the accuracy of the models against the ground truth. The datasets, however, are typically biased towards more common scenarios, such as cruising. Consequently, when metrics are averaged over the entire data, the benchmark tends to favor the models that perform better on common cases and provides little insight into the differences among the models and how they behave in rarer situations. This would limit the ability to observe the effects of design choices on particular driving scenarios.
@@
Furthermore, we propose a novel set of bias-free metrics for accurately measuring the performance of trajectory prediction models. Our metrics present a new insight into the diversity and admissibility of trajectories by defining road and kinematic compliancy for normal driving behavior, which are vital aspects of real-world driving. Our metrics help characterize models’ behavior more effectively and provide a different, yet more accurate ranking scheme than the existing metrics.
@@ II-A Trajectory Prediction
An extensive body of literature is dedicated to trajectory prediction in the context of autonomous driving [7, 8, 9, 10, 11, 12, 13, 14, 15]. Existing methods investigate various representations and approaches to learning scene context, including rasterized images [16, 17, 18], point-clouds [19] processed with convolutional neural networks, and vectorized representations processed using different architectures, such as graph neural networks [20, 21, 22, 23] or transformer architectures[24, 25, 26, 27] often combined with sophisticated fusion mechanisms [28]. To address the inherent uncertainty in predicting future trajectories, many models resort to generating multimodal trajectories in order to cover the space of possibilities. There are different approaches to achieving this goal, including the use of generative adversarial networks [29] [30], conditional variational autoencoders [31], or training sampling networks [32, 5]. Some methods also use trajectory endpoints [22] [33] or targets [34, 22, 35, 33] to model possible future intentions of vehicles to encourage more diverse predictions. Regardless of the method of choice, there is a need for an effective mechanism to evaluate the quality of generated trajectories both in terms of diversity and admissibility, i.e. compliancy to road structure or dynamical limitations.
@@ II-B1 Distance-based Metrics
II-B2 Admissibility Metrics
Admissibility metrics determine whether trajectories satisfy map constraints, e.g. road boundaries. Some of the metrics in this category include Driveable Area Compliance (DAC) and off-road rate [1], which measure the ratio of trajectories that fall within the driveable area. Both of these metrics, however, consider the entire scene as the drivable area and do not penalize the model if the predictions go outside of the valid lanes, e.g. the ones in the opposite direction. Another metric is referred to as Driveable Area Occupancy (DAO) [4], which measures the proportion of pixels occupied by the trajectories in the drivable area. Similar to other metrics, DAO is highly biased towards the length of trajectories and it does not constrain the extent of the drivable area. This means, the longer trajectories get the higher (better) the value of the metric.
@@ III-A Problem Formulation
Trajectory prediction task can be viewed as an optimization problem. At time step $t$, let consider the past trajectory of the $i$-th agent as a set of 2D coordinates in bird’s eye view over some observation horizon $L$ time steps $X_{i}=\{(x_{i},y_{i})^{t-L+1},\cdots,(x_{i},y_{i})^{t}\}$. Then, the objective is to predict future trajectories $Y_{i}=\{(x_{i},y_{i})^{t+1},\cdots,(x_{i},y_{i})^{t+T}\}$, where $T$ is the prediction horizon. In this formulation, additional contextual information, such as road lane boundaries and links are used. In the multimodal setting, prediction models output $K$ different trajectories (referred to as modes) per each vehicle. For simplicity, in the remainder of this paper, we refer to $t-L+1$ and $t+T$ as $L$ and $T$, respectively.
@@ III-B Scenario Extraction
Trajectory prediction models have different architectural designs and often behave differently when applied to the same samples. For instance, the same models may generate different distributions, and trajectories with different lengths, curvature, etc. As a result, it is essential to extract driving scenarios in a meaningful way to better reflect such differences among the models. To this end, we propose a method to extract driving scenarios based on three criteria, namely the road structure, models’ performance, and data properties.
@@ III-C2 Road Boundary Alignment Test
$\mathcal{C}=max(0,1-\frac{\Delta\Theta}{\pi}).$ (1)
@@ III-D2 AMV: Average Magnitude Variation
where $\|\vec{v^{\prime}}^{t}\|$ is the magnitude of the trajectory vector $\vec{v}=\langle\hat{x}^{t}-\hat{x}^{t-1},\hat{y}^{t}-\hat{y}^{t-1}\rangle$ starting at time step $t-1$ and ending at $t$ that is passed by kinematic the compliancy test. The magnitude variation between two trajectories is calculated once $MV_{ij}=MV_{ji}$, and then AMV is calculated by taking an average over all trajectory pairs’ magnitude variances.
@@ IV-B Accuracy, Diversity, and Admissibility
The proposed diversity metrics show an alternative ranking. Thanks to removing length bias, we can see that MMTransformer achieves the highest magnitude diversity. This new ranking suggests that TNT tends to generate longer trajectories, compared to other models, hence being favored in diversity metrics in general. This model, however, achieves a second best score in terms of angular diversity owing to target selection based on road topology.
@@
Figure 2: Performance comparison on the challenging scenarios in terms of diversity, admissibility, and minFDE (represented as the size of circles, so smaller is better).
@@ IV-C Scenario-based Analysis
Figure 3: Qualitative samples of different ATT test scores for predicted trajectories. The red lines indicate inadmissible trajectories. From left to right, the samples are generated by TNT, MMTransformer, TNT and LaneGCN.